I understand how you answer the Questions that say:
Write an equation of the line that passes through the point and has the given slope.Write the equation in slope intercept form.
and they give you
(-1,0),m=3
However im stuck on the next step where they say:
Write an equation in slope intercept form of the line that passes through the points.
and they give you
(3,-2),(5,4)
Im really confused and i forgot how to do it. i have math in the mornings and when i get home i simply dont remember. If you could help that would be great. thanks,
Kirsten archer.
(They Being McDougal littell 2004 edition.)

Found 2 solutions by solver91311, edjones:
Answer by solver91311(24713)



You need the two-point form of the equation of a line:



Where and are the coordinates of the given points. Just plug in the values, do the indicated arithmetic, and rearrange the equation to slope-intercept form.

This is good deal easier to remember if you just remember your point-slope form:



and then the slope formula:



Notice that the two-point form is exactly the same as the point-slope form with the formula for substituted in place of

John


Answer by edjones(8007)

m=3 (-1, 0)
y-y[1]=m(x-x[1])
y-0=3(x-(-1))
y=3x+3 slope intercept form.
.
(3, -2), (5, 4)
(y-y[1])/(x-x[1])=m
(-2-4)/(3-5)
=-6/-2
=3=m
y-4=3(x-5)
y-4=3x-15
y=3x-11
.
Ed